# Improving the fact-checking performance of language models by relying on their entailment ability

### v1

@@ 4 Results and Discussion:
In this section, we have reported the results from all of our experiments. While we used macro-precision, macro-recall and macro-F1 scores to evaluate the veracity, ROUGE, BLEU and BERT-score, along with subjective evaluation by VLLMs, were made to evaluate the explanations. A detailed description of evaluation metrics is reported in subsection C.6.4 in the appendix.
Dataset ($\xrightarrow{}$) LIAR-RAW RAW-FC
TBE-1 TBE-1
Method ($\downarrow$) MP MR MF1 MP MR MF1
@@
     -Mistral 0.44 0.29 0.27 0.69 0.65 0.65
(±0.02) (±0.01) (±0.01) (±0.01) (±0.00) (±0.01)
     -Llama 0.34 0.30 0.30 0.68 0.64 0.65
(±0.01) (±0.02) (±0.00) (±0.01) (±0.02) (±0.02)
     -Gemma 0.27 0.25 0.23 0.60 0.58 0.57
(±0.02) (±0.03) (±0.04) (±0.05) (±0.04) (±0.06)
     -Qwen 0.37 0.29 0.29 0.67 0.66 0.66
(±0.02) (±0.02) (±0.04) (±0.03) (±0.03) (±0.03)
     -Falcon 0.39 0.28 0.26 0.59 0.58 0.54
(±0.01) (±0.01) (±0.01) (±0.03) (±0.05) (±0.06)
@@
     -Mistral 0.40 0.27 0.25 0.53 0.56 0.55
(±0.03) (±0.02) (±0.01) (±0.05) (±0.03) (±0.02)
     -Llama 0.34 0.29 0.29 0.67 0.64 0.65
(±0.01) (±0.01) (±0.01) (±0.01) (±0.01) (±0.01)
     -Gemma 0.27 0.23 0.22 0.61 0.57 0.57
(±0.02) (±0.02) (±0.03) (±0.03) (±0.02) (±0.03)
     -Qwen 0.36 0.29 0.29 0.70 0.65 0.65
(±0.01) (±0.01) (±0.02) (±0.02) (±0.02) (±0.02)
     -Falcon 0.37 0.30 0.29 0.64 0.62 0.63
(±0.02) (±0.01) (±0.02) (±0.03) (±0.03) (±0.03)
Table 3: Performance of LoRA and LoRA+ models on TBE-1 (LIAR-RAW and RAW-FC datasets).
4.1 Observations from TBE-1:
We reported the performance of various models for TBE-1, i.e. training with raw evidence sentences as input in Table 3. We found that LoRA adapter trained with Llama VLLM resulted in the highest F1 score of 0.30 for the LIAR-RAW dataset. While this performance is comparable to the performance of baseline models L-Defense ($F1$: 0.31) and FactLLaMA ($F1$: 0.30), it does not surpass the performance of other baselines, i.e. HiSS ($F1$: 0.37) and RAFTS ($F1$: 0.42). However, for the RAWFC dataset, we found that many models, such as Mistral with LoRA ($F1$: 0.65, $\sim\textbf{6.56}\%\uparrow$), Llama with LoRA ($F1$: 0.65, $\sim\textbf{6.56}\%\uparrow$), Qwen with LoRA ($F1$: 0.66, $\sim\textbf{8.20}\%\uparrow$), Llama with LoRA+ ($F1$: 0.65, $\sim\textbf{6.56}\%\uparrow$), Qwen with LoRA+ ($F1$: 0.65, $\sim\textbf{6.56}\%\uparrow$) and Falcon with LoRA+ ($F1$: 0.63, $\sim\textbf{3.28}\%\uparrow$), outperform the best baseline performance ($F1$: 0.61). Here, $\sim x\%\uparrow$ denotes the relative percentage improvement compared to the best baseline score. Qwen trained with LoRA resulted in the overall highest F1-score (0.66, $\sim\textbf{8.20}\%\uparrow$). We have reported some additional observations in the appendix section D.1.
Dataset ($\xrightarrow{}$) LIAR-RAW RAW-FC
Method ($\downarrow$) TBE-2 TBE-3 TBE-2 TBE-3
MP MR MF1 MP MR MF1 MP MR MF1 MP MR MF1
Fine-tuning
     -RoBERTa-LMistral 0.28 0.26 0.26 0.48 0.47 0.47 0.51 0.50 0.50 0.83 0.82 0.83
(±0.01) (±0.01) (±0.01) (±0.01) (±0.00) (±0.01) (±0.01) (±0.01) (±0.01) (±0.00) (±0.01) (±0.01)
     -RoBERTa-LLlama 0.27 0.28 0.25 0.53 0.53 0.52 0.50 0.50 0.49 0.88 0.88 0.88
(±0.02) (±0.01) (±0.02) (±0.01) (±0.01) (±0.01) (±0.01) (±0.01) (±0.01) (±0.01) (±0.01) (±0.01)
     -RoBERTa-LGemma 0.28 0.27 0.27 0.49 0.50 0.48 0.51 0.51 0.50 0.50 0.49 0.49
(±0.02) (±0.01) (±0.01) (±0.01) (±0.02) (±0.02) (±0.04) (±0.03) (±0.04) (±0.02) (±0.01) (±0.02)
     -RoBERTa-LQwen 0.30 0.28 0.28 0.48 0.47 0.46 0.51 0.49 0.48 0.73 0.71 0.71
(±0.01) (±0.02) (±0.01) (±0.01) (±0.02) (±0.01) (±0.01) (±0.01) (±0.02) (±0.02) (±0.02) (±0.02)
     -RoBERTa-LFalcon 0.29 0.28 0.27 0.49 0.43 0.44 0.50 0.48 0.48 0.65 0.64 0.64
(±0.02) (±0.02) (±0.01) (±0.01) (±0.02) (±0.01) (±0.00) (±0.01) (±0.01) (±0.02) (±0.02) (±0.02)
     -XLNet-LMistral 0.29 0.28 0.28 0.49 0.47 0.47 0.62 0.61 0.61 0.83 0.82 0.82
(±0.02) (±0.01) (±0.01) (±0.00) (±0.01) (±0.01) (±0.02) (±0.01) (±0.02) (±0.01) (±0.01) (±0.01)
     -XLNet-LLlama 0.31 0.29 0.29 0.55 0.54 0.54 0.63 0.63 0.62 0.88 0.88 0.87
(±0.02) (±0.01) (±0.01) (±0.01) (±0.02) (±0.01) (±0.03) (±0.03) (±0.03) (±0.01) (±0.01) (±0.01)
     -XLNet-LGemma 0.26 0.26 0.25 0.47 0.43 0.42 0.51 0.50 0.50 0.47 0.47 0.46
(±0.02) (±0.02) (±0.02) (±0.04) (±0.05) (±0.09) (±0.01) (±0.01) (±0.01) (±0.02) (±0.02) (±0.02)
     -XLNet-LQwen 0.29 0.29 0.28 0.50 0.48 0.48 0.60 0.58 0.58 0.70 0.70 0.70
(±0.02) (±0.02) (±0.02) (±0.01) (±0.01) (±0.01) (±0.02) (±0.01) (±0.01) (±0.03) (±0.03) (±0.04)
     -XLNet-LFalcon 0.26 0.26 0.24 0.45 0.44 0.44 0.61 0.60 0.60 0.76 0.75 0.74
(±0.03) (±0.01) (±0.03) (±0.01) (±0.02) (±0.01) (±0.02) (±0.01) (±0.01) (±0.01) (±0.01) (±0.01)
LoRA
     -Mistral 0.36 0.32 0.32 0.35 0.34 0.29 0.61 0.60 0.60 0.69 0.59 0.58
(±0.01) (±0.01) (±0.01) (±0.04) (±0.03) (±0.03) (±0.04) (±0.04) (±0.04) (±0.05) (±0.05) (±0.05)
     -Llama 0.31 0.27 0.25 0.36 0.32 0.30 0.48 0.48 0.46 0.63 0.62 0.61
(±0.04) (±0.03) (±0.04) (±0.03) (±0.03) (±0.04) (±0.06) (±0.03) (±0.02) (±0.03) (±0.03) (±0.03)
     -Gemma 0.21 0.22 0.18 0.26 0.23 0.20 0.47 0.45 0.40 0.36 0.34 0.30
(±0.01) (±0.01) (±0.01) (±0.03) (±0.03) (±0.02) (±0.03) (±0.03) (±0.05) (±0.02) (±0.01) (±0.02)
     -Qwen 0.32 0.27 0.21 0.40 0.36 0.32 0.56 0.54 0.53 0.48 0.46 0.46
(±0.09) (±0.02) (±0.02) (±0.07) (±0.07) (±0.05) (±0.07) (±0.0) (±0.03) (±0.03) (±0.03) (±0.03)
     -Falcon 0.28 0.26 0.23 0.27 0.24 0.21 0.56 0.57 0.53 0.43 0.0.40 0.41
(±0.01) (±0.00) (±0.01) (±0.03) (±0.03) (±0.02) (±0.03) (±0.01) (±0.00) (±0.01) (±0.03) (±0.03)
LoRA+
     -Mistral 0.33 0.30 0.30 0.30 0.32 0.27 0.66 0.60 0.60 0.52 0.48 0.46
(±0.01) (±0.01) (±0.01) (±0.07) (±0.02) (±0.03) (±0.05) (±0.05) (±0.05) (±0.06) (±0.04) (±0.04)
     -Llama 0.27 0.24 0.23 0.49 0.50 0.49 0.73 0.71 0.71 0.84 0.82 0.82
(±0.01) (±0.02) (±0.02) (±0.03) (±0.02) (±0.02) (±0.02) (±0.02) (±0.02) (±0.01) (±0.03) (±0.03)
     -Gemma 0.23 0.21 0.20 0.34 0.32 029 0.53 0.51 0.51 0.52 0.48 0.46
(±0.03) (±0.02) (±0.02) (±0.04) (±0.05) (±0.06) (±0.02) (±0.02) (±0.04) (±0.03) (±0.04) (±0.01)
     -Qwen 0.27 0.27 0.24 0.33 0.34 0.29 0.57 0.48 0.43 0.52 0.49 0.47
(±0.04) (±0.03) (±0.04) (±0.03) (±0.03) (±0.03) (±0.07) (±0.03) (±0.01) (±0.06) (±0.09) (±0.09))
     -Falcon 0.33 0.28 0.28 0.39 0.39 0.36 0.55 0.56 0.53 0.58 0.52 0.50
(±0.00) (±0.01) (±0.02) (±0.02) (±0.03) (±0.04) (±0.03) (±0.02) (±0.03) (±0.01) (±0.02) (±0.03)
@@ 4.1 Observations from TBE-1:
Table 4: Performance of claim veracity prediction using gold evidences. Green and Blue indicates best and second-best performance, respectively.
4.2 Observations from TBE-2:
We reported the performance of various models for TBE-2, i.e. training with claim-evidence understanding as input in table 4. We found that Mistral trained with LoRA resulted in the highest macro-F1 (0.32) for the LIAR-RAW dataset. While its performance surpasses the performance of baselines FactLLaMA ($MF1:\textbf{0.30}$) and L-Defense ($MF1:\textbf{0.31}$), it is still behind HiSS ($MF1:\textbf{0.37}$) and RAFTS ($MF1:\textbf{0.42}$). However, for the RAW-FC dataset, we observed that many models, such as XLNet fine-tuned on Llama based understandings ($MF1:$ 0.62, $\sim\textbf{1.64}\%\uparrow$), and Llama trained (with Llama understandings) with LoRA+ ($MF1:\textbf{0.71}$, $\sim\textbf{16.39}\%\uparrow$) outperformed the best reported macro-F1 score by the baselines ($MF1:\textbf{0.61}$). Llama trained (with Llama understandings) with LoRA+ adapter resulted in the highest overall performance ($MF1:$ 0.71, $\sim\textbf{16.39}\%\uparrow$). We have reported some additional observations in the section D.2 of the appendix due to space constraints.
4.3 Observations from TBE-3:
We reported the performance of various models for TBE-3, i.e. training with entailed justifications generated by VLLMs as input in table 4. We found that XLNet fine-tuned on Llama based entailed justification resulted in the highest macro-F1 (0.54, $\sim\textbf{28.57}\%\uparrow$) for the LIAR-RAW dataset. Many models, such as RoBERTa fine-tuned with Mistral ($MF1:\textbf{0.47}$, $\sim\textbf{16.39}\%\uparrow$), Llama ($MF1:\textbf{0.52}$, $\sim\textbf{23.81}\%\uparrow$), Gemma ($MF1:\textbf{0.48}$, $\sim\textbf{14.29}\%\uparrow$), Qwen ($MF1:\textbf{0.46}$, $\sim\textbf{9.52}\%\uparrow$), and Falcon ($MF1:\textbf{0.44}$, $\sim\textbf{4.76}\%\uparrow$) based entailed justifications, XLNet fine-tuned with Mistral ($MF1:\textbf{0.47}$, $\sim\textbf{16.39}\%\uparrow$), Llama ($MF1:\textbf{0.54}$, $\sim\textbf{28.57}\%\uparrow$), Qwen ($MF1:\textbf{0.48}$, $\sim\textbf{14.29}\%\uparrow$), and Falcon ($MF1:\textbf{0.44}$, $\sim\textbf{4.76}\%\uparrow$) based entailed justifications, and Llama trained (with Llama justifications) with LoRA+ adapter ($MF1:\textbf{0.49}$, $\sim\textbf{16.67}\%\uparrow$) surpassed the best reported macro-F1 score of baselines ($MF1:\textbf{0.42}$). In a similar manner, for the RAW-FC dataset, we observed that many models, such as such as RoBERTa fine-tuned with Mistral ($MF1:\textbf{0.83}$, $\sim\textbf{36.07}\%\uparrow$), Llama ($MF1:\textbf{0.88}$, $\sim\textbf{44.26}\%\uparrow$), Qwen ($MF1:\textbf{0.71}$, $\sim\textbf{16.39}\%\uparrow$), and Falcon ($MF1:\textbf{0.64}$, $\sim\textbf{4.91}\%\uparrow$) based entailed justifications, XLNet fine-tuned with Mistral ($MF1:\textbf{0.82}$, $\sim\textbf{34.42}\%\uparrow$), Llama ($MF1:\textbf{0.87}$, $\sim\textbf{42.62}\%\uparrow$), Qwen ($MF1:\textbf{0.70}$, $\sim\textbf{14.75}\%\uparrow$), and Falcon ($MF1:\textbf{0.74}$, $\sim\textbf{21.31}\%\uparrow$) based entailed justifications, and Llama trained (with Llama justifications) with LoRA+ adapter ($MF1:\textbf{0.83}$, $\sim\textbf{36.07}\%\uparrow$) outperformed the best reported macro-F1 score by the baselines ($MF1:\textbf{0.61}$). RoBERTa fine-tuned with Mistral based entailed justification achieved the highest overall performance ($MF1:$ 0.88, $\sim\textbf{44.26}\%\uparrow$). We have reported some additional observations in the Appendix section D.3.
Dataset ($\xrightarrow{}$) LIAR-RAW RAW-FC
@@ 4.3 Observations from TBE-3:
     -Mistral 0.24 0.25 0.22 0.22 0.23 0.20 0.40 0.23 0.21 0.30 0.17 0.14 0.54 0.54 0.53 0.58 0.58 0.58 0.50 0.52 0.45 0.45 0.46 0.43
     -Llama 0.24 0.23 0.20 0.30 0.23 0.22 0.26 0.22 0.21 0.22 0.13 0.13 0.56 0.56 0.54 0.62 0.63 0.62 0.45 0.47 0.49 0.42 0.39 0.35
     -Gemma 0.24 0.21 0.19 0.16 0.16 0.13 0.27 0.19 0.16 0.09 0.16 0.11 0.40 0.40 0.40 0.41 0.38 0.38 0.45 0.41 0.40 0.27 0.31 0.24
     -Qwen 0.27 0.23 0.20 0.25 0.23 0.20 0.24 0.22 0.21 0.13 0.16 0.13 0.61 0.58 0.59 0.58 0.57 0.57 0.57 0.54 0.52 0.50 0.46 0.43
     -Falcon 0.24 0.23 0.20 0.22 0.22 0.20 0.24 0.23 0.21 0.16 0.17 0.14 0.56 0.57 0.54 0.60 0.59 0.57 0.60 0.52 0.48 0.40 0.38 0.37

### v2

@@ 4 Results and Discussion:
In this section, we have reported the results of our experiments. We used macro-precision ($MP$), macro-recall ($MR$) and macro-F1 ($MF1$) scores to evaluate the veracity prediction. Similarly, we used ROUGE, BLEU, BERT-score and have done subjective evaluation by VLLMs to evaluate model explainability. A detailed description of evaluation metrics is reported in subsection C.6.4 of the appendix.
Dataset Method Mistral Llama Gemma Qwen Falcon
LIAR-RAW LoRA 0.27 (± 0.01) 0.30 (± 0.00) 0.23 (± 0.04) 0.29 (± 0.04) 0.26 (± 0.01)
LoRA+ 0.25 (± 0.01) 0.29 (± 0.01) 0.22 (± 0.03) 0.29 (± 0.02) 0.29 (± 0.02)
RAW-FC LoRA 0.65 (± 0.01) 0.65 (± 0.02) 0.57 (± 0.06) 0.66 (± 0.03) 0.54 (± 0.06)
LoRA+ 0.55 (± 0.02) 0.65 (± 0.01) 0.57 (± 0.03) 0.65 (± 0.02) 0.63 (± 0.03)
Table 3: Performance of models under TBE-1 for LIAR-RAW and RAW-FC datasets in terms of $MF1$. We have also reported variation across three random seeds (±x). Green and Blue indicate best and second-best scores, respectively.
4.1 Observations for veracity prediction:
We reported the $MF1$ of various models under TBE-1, TBE-2 - TBE-3 and IBEs in Table 3, Table 4, and Table 5 respectively. A comparison of best performing models across all experiments are illustrated in Figure 22. Our observations are as follows,
  • For LIAR-RAW, none of the macro-F1 reported by IBE models could surpass the best baseline performance. In contrast, for RAW-FC, Llama achieved the highest performance in IBE-2, surpassing the highest baseline performance.
  • In TBE-1, for LIAR-RAW, the highest reported $MF1$ does not surpass that of baselines. However, for RAWFC, we found that many models, such as (i) Mistral ($\sim\textbf{6.56}\%\uparrow$), Llama ($\sim\textbf{6.56}\%\uparrow$) and Qwen ($\sim\textbf{8.20}\%\uparrow$) trained with LoRA, and (ii) Llama ($\sim\textbf{6.56}\%\uparrow$), Qwen ($\sim\textbf{6.56}\%\uparrow$) and Falcon ($\sim\textbf{3.28}\%\uparrow$) trained with LoRA+, outperform the best baseline performance ($F1$: 0.61).
  • In TBE-2, for LIAR-RAW, no model surpassed the best $MF1$ reported by the baselines. However, for RAW-FC, we observed that many models, such as (i) XLNet fine-tuned on Llama understandings ($\sim\textbf{1.64}\%\uparrow$), and (ii) Llama trained with LoRA+ based on Llama understandings ($\sim\textbf{16.39}\%\uparrow$) outperformed the best reported baseline $MF1$.
  • Many models in TBE-3, such as (i) RoBERTa fine-tuned with Mistral ($\sim\textbf{16.39}\%\uparrow$), Llama ($\sim\textbf{23.81}\%\uparrow$), Gemma ($\sim\textbf{14.29}\%\uparrow$), Qwen ($\sim\textbf{9.52}\%\uparrow$), and Falcon ($\sim\textbf{4.76}\%\uparrow$) based entailed justifications, (ii) XLNet fine-tuned with Mistral ($\sim\textbf{16.39}\%\uparrow$), Llama ($\sim\textbf{28.57}\%\uparrow$), Qwen ($\sim\textbf{14.29}\%\uparrow$), and Falcon ($\sim\textbf{4.76}\%\uparrow$) based entailed justifications, and (iii) Llama trained with Llama justifications and LoRA+ adapter ($\sim\textbf{16.67}\%\uparrow$) surpassed the best reported macro-F1 score of baselines ($MF1:\textbf{0.42}$). Similarly, for RAW-FC, we observed that (i) RoBERTa fine-tuned with Mistral ($\sim\textbf{36.07}\%\uparrow$), Llama ($\sim\textbf{44.26}\%\uparrow$), Qwen ($\sim\textbf{16.39}\%\uparrow$), and Falcon ($\sim\textbf{4.91}\%\uparrow$) based entailed justifications, (ii) XLNet fine-tuned with Mistral ($\sim\textbf{34.42}\%\uparrow$), Llama ($\sim\textbf{42.62}\%\uparrow$), Qwen ($\sim\textbf{14.75}\%\uparrow$), and Falcon ($\sim\textbf{21.31}\%\uparrow$) based entailed justifications, and (iii) Llama trained (with Llama justifications) with LoRA+ adapter ($\sim\textbf{36.07}\%\uparrow$) outperformed the best reported macro-F1 score by the baselines ($MF1:\textbf{0.61}$). Models from TBE-3, exhibit highest overall performance. Figure 22 illustrates the same.
Some additional observations based on $MP$ and $MR$ are reported in the appendix (section D).
[…]
Method ($\downarrow$) TBE-2 TBE-3 TBE-2 TBE-3
Fine-tuning
     -RoBERTa-LMistral 0.26 0.47 0.50 0.83
(±0.01) (±0.01) (±0.01) (±0.01)
     -RoBERTa-LLlama 0.25 0.52 0.49 0.88
(±0.02) (±0.01) (±0.01) (±0.01)
     -RoBERTa-LGemma 0.27 0.48 0.50 0.49
(±0.01) (±0.02) (±0.04) (±0.02)
     -RoBERTa-LQwen 0.28 0.46 0.48 0.71
(±0.01) (±0.01) (±0.02) (±0.02)
     -RoBERTa-LFalcon 0.27 0.44 0.48 0.64
(±0.01) (±0.01) (±0.01) (±0.02)
     -XLNet-LMistral 0.28 0.47 0.61 0.82
(±0.01) (±0.01) (±0.02) (±0.01)
     -XLNet-LLlama 0.29 0.54 0.62 0.87
(±0.01) (±0.01) (±0.03) (±0.01)
     -XLNet-LGemma 0.25 0.42 0.50 0.46
(±0.02) (±0.09) (±0.01) (±0.02)
     -XLNet-LQwen 0.28 0.48 0.58 0.70
(±0.02) (±0.01) (±0.01) (±0.04)
     -XLNet-LFalcon 0.24 0.44 0.60 0.74
(±0.03) (±0.01) (±0.01) (±0.01)
@@ 4.1 Observations for veracity prediction:
     -Mistral 0.32 0.29 0.60 0.58
(±0.01) (±0.03) (±0.04) (±0.05)
     -Llama 0.25 0.30 0.46 0.61
(±0.04) (±0.04) (±0.02) (±0.03)
     -Gemma 0.18 0.20 0.40 0.30
(±0.01) (±0.02) (±0.05) (±0.02)
     -Qwen 0.21 0.32 0.53 0.46
(±0.02) (±0.05) (±0.03) (±0.03)
     -Falcon 0.23 0.21 0.53 0.41
(±0.01) (±0.02) (±0.00) (±0.03)
@@
     -Mistral 0.30 0.27 0.60 0.46
(±0.01) (±0.03) (±0.05) (±0.04)
     -Llama 0.23 0.49 0.71 0.82
(±0.02) (±0.02) (±0.02) (±0.03)
     -Gemma 0.20 0.29 0.51 0.46
(±0.02) (±0.06) (±0.04) (±0.01)
     -Qwen 0.24 0.29 0.43 0.47
(±0.04) (±0.03) (±0.01) (±0.09)
     -Falcon 0.28 0.36 0.53 0.50
(±0.02) (±0.04) (±0.03) (±0.03)
@@
Table 4: Performance of claim veracity prediction using gold evidences in terms of macro F1(MF) score. Green and Blue indicate best and second-best performance, respectively.
[…]
@@
     -Mistral 0.22 0.20 0.21 0.14 0.53 0.58 0.45 0.43
     -Llama 0.20 0.22 0.21 0.13 0.54 0.62 0.49 0.35
     -Gemma 0.19 0.13 0.16 0.11 0.40 0.38 0.40 0.24
     -Qwen 0.20 0.20 0.21 0.13 0.59 0.57 0.52 0.43
     -Falcon 0.20 0.20 0.21 0.14 0.54 0.57 0.48 0.37